OpenStudy (anonymous):

750 is what percent of 900 (Round to nearest tenth of a percent).

OpenStudy (anonymous):

80%

OpenStudy (anonymous):

750/900 * 100 = 83.3333

OpenStudy (anonymous):

shld be 83.3% nearest tenth

OpenStudy (radar):

Let x = the decimal equivalent of the percent x*900=750 x=750/900=.8333 convert decimal to per cent by multiplying by 100=83.33%